Circular No. 8785

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 YC (CATALINA-CHRISTENSEN) Tabulated below are the initial observations of a comet that was discov-

ered almost simultaneously via the Catalina Sky Survey (with the observer, A. R. Gibbs, reporting the object as apparently asteroidal on the 0.68-m Schmidt telescope exposures) and the Mount Lemmon Survey (whose observer, E. J. Christensen, reported a 6′′ coma and 10′′ tail in p.a. 270^{o} on four 30-s co-added discovery images taken with the 1.5-m reflector). Images taken in fair-to-poor seeing at Mt. Lemmon by E. C. Beshore on Dec. 20.5 show a moderately condensed coma ∼ 6′′ in diameter with a 10′′ tail in p.a. 250^{o}. When K. Smalley designated the object as 2006 YC via MPEC 2006-Y15 on Dec. 17, he unfortunately was unaware of Christensen’s report (submitted to the Central Bureau and to other Minor Planet Center staff members) of the object’s cometary appearance.

2006 UT α_{2000} δ_{2000} Mag. Observer
Dec. 16.41622 10^{h}31^{m}38.^{s}56 −4^{o}08′53′′.4 20.0 Gibbs
16.48765 10 31 38.94 −4 08 33.2 19.4 Christensen

Additional astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2006-Y26.

T = 2007 Sept.25.985 TT  ω = 29.^{o}255
Ω = 152.881 }2000.0
q = 4.23768 AU  i = 73.988

V2362 CYGNI D. K. Lynch, R. W. Russell, D. Kim, The Aerospace Corporation; and

M. L. Sitko and S. Brafford, University of Cincinnati and Space Science Institute, report that 3- to 13-μm spectroscopy of V2362 Cyg (cf. IAUC 8731) on Dec. 12.25 UT using BASS on the Infrared Telescope Facility showed a smooth stellar-like continuum decreasing monotonically toward longer wavelengths between 3 and 13 μm with no evidence of thermal emission from dust. Infrared magnitudes at that time were L = 5.0 ± 0.1, M = 4.5 ± 0.2, N = 4.0 ± 0.2.

Visual magnitude estimates: Nov. 15.736 UT, 11.5 (M. Lehky, Hradec Kralove, Czech Republic); 18.731, 10.9 (A. Diepvens, Balen, Belgium); Dec. 10.836, 12.0 (J. Carvajal, Madrid, Spain).

2006 December 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green